# On the maximal correlation of some stochastic processes

### v2

@@ 4 Lévy processes and stable distributions
Since $X_{1}$ is stable, there exists $(c_{n},d_{n})\in\mathbb{R}$ such that $(S_{n}-c_{n},T_{n}-d_{n})$ has the same distribution as $n^{1/\alpha}(X_{1},Y_{1})$. Define
@@
Then, $R(X^{(n)},Y^{(n)})=R(S,T)=R(X_{1},Y_{1})$. Moreover, by [Kal21, Theorem 23.14], the process $(X^{(n)}_{t},Y^{(n)}_{t})_{t\geq 0}$ converges towards $(X_{t},Y_{t})_{t\geq 0}$ in the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$. Hence, by Lemma 2.3, we have that
@@
Suppose that $(X,Y)$ is a stable random vector. Let $(\widetilde{X},\widetilde{Y})$ be in the domain of contraction of $(X,Y)$. Then, we have that
@@ 4.3 Examples
For $\alpha=2$, $X$ and $Z$ must be Gaussian and the result goes back to Lancaster [Lan57]. The main contribution is the case that $20<\alpha<$ 2, $\lambda\geq 0$ and $X$ is $\alpha$-stable. The result for $\lambda<0$ and symmetric $\alpha$-stable random variables can be easily deduced from (26) by taking $-Z$ instead of $Z$. The restriction to positive $\lambda$ in (26) has a reason: In general, the expression for $\lambda<0$ is different. Indeed, we have the following result:
@@
Let $X$ and $Z$ be independent copies of $\alpha$-stable random variables with $20<\alpha<$ 2. Then, the Lévy measure $\nu_{X}(x)\mathrm{d}x$ of $X$ has the following form:
@@
By using characteristic functions and independence between $X$ and $Z$, we find that $(X,Y)=(X,X+\lambda Z)$ is $\alpha$-stable random vector with Lévy measure $\nu(\mathrm{d}x,\mathrm{d}y)$, where
@@
Recall the notations in the statement of Theorem 1.3. Then, in our case, we have that
@@
$R(X,Y)$ $=\left\|\begin{pmatrix}\sqrt{\frac{c_{+}}{c_{+}+c_{-}|\lambda|^{\alpha}}}&0\\ 0&\sqrt{\frac{c_{-}}{c_{-}+c_{+}|\lambda|^{\alpha}}}\end{pmatrix}\right\|_{2}$
@@
Secondly, we consider several compound Poisson processes.
@@
$\nu(\{(x,y)\in\mathbb{R}^{2}\setminus\{0\}:x=0\text{ or }y=0\})=0.$
Let $(X_{t},Y_{t})$ be the corresponding Lévy processes. Then, we have that
@@
The reason is that the number of jumps $N$ of $(X_{t})_{t\geq 0}$ up to time $1$ is equal to that of $(Y_{t})_{t\geq 0}$. Hence, $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq R(N,N)=1$. So, for generic compound Poisson processes, the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ is equal to $1$.
Finally, we show that the maximal correlation coefficient $R(X_{1},Y_{1})$ could be strictly smaller than $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$.
@@
Let $(B_{t})_{t\geq 0}$ be a (one dimensional) Brownian motion. Let $(N_{t})_{t\geq 0}$ be a Poisson process of rate $1$ which is independent of $(B_{t})_{t\geq 0}$. Let $X_{t}=B_{t}+N_{t}$. Then, $(X_{t},N_{t})_{t\geq 0}$ is a two dimensional Lévy process. By Theorem 1.2, we have that $R((X_{t})_{t\geq 0},(N_{t})_{t\geq 0})\geq\mathrm{Op}(\nu)=R((N_{t})_{t\geq 0},(N_{t})_{t\geq 0})=1$. Since the maximal correlation coefficients lie in the interval $[0,1]$, we have that $R((X_{t})_{t\geq 0},(N_{t})_{t\geq 0})=1$. However, $R(X_{1},N_{1})<1$. We will show this by contradiction. Assume that $R(X_{1},N_{1})=1$. Let $f(x,n)\,\mathrm{d}x=P(X_{1}\in(x,x+\mathrm{d}x),N_{1}=n)$, i.e.
@@
where $f_{N_{1}|X_{1}}(n|x)=P(N_{1}=n|X_{1}=x)$ and $f_{X_{1}|N_{1}}(x|n)$ is the conditional density of $X_{1}$ given $N_{1}=n$. Here, $C$ is called mean square contingency of $X_{1}$ and $N_{1}$. We will show that $C<\infty$ as follows: Note that
@@
Then, there exists $c>0$ such that for $x\geq e$, we have that
@@
where $\lceil x\rceil$ is the least integer that is greater than or equal to $x$. Then,
$f_{X_{1}}(x)=\sum_{n=0}^{\infty}f(x,n)\geq\max_{n\geq 0}f(x,n)=g(x)$
and for $x\geq e$,
@@
Combining (29), (32) and (33), we get (31). Hence, the mean square contingency $C$ is finite. By [R5́9, Theorem 2], there exists non-degenerate $\varphi(X_{1})$ and $\psi(N_{1})$ such that $\rho(\varphi(X_{1}),\psi(N_{1}))=R(X_{1},N_{1})=1$. Therefore, there exist $c_{1}\neq 0,c_{2}\neq 0$ and $d\in\mathbb{R}$ such that $P(c_{1}\varphi(X_{1})+c_{2}\psi(N_{1})=d)=1$. However, given the value of $N_{1}$, $X_{1}$ could take any value in $\mathbb{R}$. If $P(c_{1}\varphi(X_{1})+c_{2}\psi(N_{1})=d)=1$, then $\varphi$ is constant almost everywhere, which contradicts with the non-degeneracy of $\varphi(X_{1})$. Finally, by contradiction, we prove that $R(X_{1},N_{1})<1$.
For compound Poisson processes, it is still possible that $R(X_{1},Y_{1})<R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$, see the following example:
@@
Let $(M_{t})_{t\geq 0}$ and $(N_{t})_{t\geq 0}$ be two independent Poisson processes of rate $1$. Let $X_{t}=M_{t}-N_{t}$ and $Y_{t}=M_{t}$. Then, $(X_{t},Y_{t})_{t\geq 0}$ is a two dimensional compound Poisson process. We have that $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq\mathrm{Op}(\nu)=1$. Since the maximal correlation coefficients is at most one, we have that $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})=1$. However, we find that $R(X_{1},Y_{1})<1$ by numerical methods although we don’t have a theoretical proof at present. Indeed, let $X^{(n)}_{1}=\min(\max(X_{1},-n),n)$ and $Y^{(n)}_{1}=\min(\max(Y_{1},-n),n)$. Since $X^{(n)}_{1}$ is a measurable function of $X_{1}$ and $Y^{(n)}_{1}$ is a measurable function of $Y_{1}$, we have that $R(X^{(n)}_{1},Y^{(n)}_{1})\leq R(X_{1},Y_{1})$. On the other hand, since $(X^{(n)}_{1},Y^{(n)}_{1})$ converges to $(X_{1},Y_{1})$ in distribution, by Lemma 2.3, we have that $R(X,Y)\leq\liminf_{n\to\infty}R(X^{(n)}_{1},Y^{(n)}_{1})$. Therefore, we must have
@@
Since $(X^{(n)}_{1},Y^{(n)}_{1})$ takes values in a finite set, the maximal correlation $R(X^{(n)}_{1},Y^{(n)}_{1})$ is given by certain eigenvalue of a finite matrix, which could be found by numerical methods. By numerical calculation, we found that $R(X,Y)=\lim_{n\to\infty}R(X^{(n)}_{1},Y^{(n)}_{1})$ is close to $0.6925$.
5 Analog and generalization of DKS inequality
5.1 Proof of Proposition 1.4
We prove Proposition 1.4 in this subsection.
Recall the definition of $(X_{1},X_{2},\ldots,X_{n})$ and $(Y_{1},Y_{2},\ldots,Y_{n})$.
Firstly, consider the case $n=1$. Let $X$ be a non-degenerate random variable. Let $B$ is a Bernoulli random variable independent of $X$. Assume that $P(B=1)=1-P(B=0)=p$. Define
@@ 5 Analog and generalization of DKS inequality
The maximum correlation coefficient between $X$ and $Y$ is equal to $\sqrt{p}$.
@@
To calculate the maximum correlation coefficient between $X$ and $Y$, we take two functions $f$ and $g$ such that $E[f(X)]=0$, $\operatorname{Var}[f(X)]<\infty$, $E[g(Y)]=0$ and $\operatorname{Var}[g(Y)]<\infty$. Then, we have that
$\operatorname{Var}[g(Y)]=E[(g(Y))^{2}]=E[E[(g(Y))^{2}|B]]=pE[(g(X))^{2}]+(1-p)(g(\partial))^{2}.$ (34)
@@
where we use Cauchy-Schwarz inequality $(E[f(X)g(X)])^{2}\leq E[(f(X))^{2}]E[(g(X))^{2}]$ in the last step. Thus, we have proved that $R(X,Y)\leq\sqrt{p}$.
Finally, there exists $g$ such that $g(\partial)=0$, $E[g(X)|X\neq\partial]=0$ and $0<E[(g(X))^{2}|X\neq\partial]<\infty$. Take $f=g$. Then, $E[f(X)]=0$ and $0<\operatorname{Var}[f(X)]=E[(g(X))^{2}]<\infty$. Moreover, $E[g(Y)]=pE[g(X)]=0$ and $0<\operatorname{Var}[g(Y)]<\infty$ by (34). Then, by (5.1), we have that
$R(X,Y)\geq\rho(f(X),g(Y))=\sqrt{p}$
@@
By Lemma 5.1, we have that
$R(X_{i},Y_{i})=\sqrt{p_{i}}.$
Since $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})$ are independent, by Csáki-Fischer identity, we have
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))$ $=\max\{R(X_{i},Y_{i}):i=1,2,\ldots,n\}$
$=\sqrt{\max(p_{1},p_{2},\ldots,p_{n})}.$
@@
In the above arguments, we need the condition that $\partial$ is out of $F$ to ensure that the distribution of $X$ conditionally on $X\neq\partial$ is non-degenerate, which guarantee the existence of $g$ such that $g(\partial)=0$, $E[g(X)|X\neq\partial]=0$ and $0<E[(g(X))^{2}|X\neq\partial]<\infty$. It is possible to allow $P(X_{i}=\partial)>0$ for some $i$. However, in such cases, we need to assume that the distribution of $X_{i}$ conditionally on $X_{i}\neq\partial$ is non-degenerate. Otherwise, the formula is no longer correct.
@@
We prove Theorem 1.6 in this subsection. The key ingredient is the analysis of variance (ANOVA) decomposition developed in [MB07, Appendix I]. For the convenience of readers, we briefly explain the ANOVA decomposition without proofs. Let $X_{1},X_{2},\ldots,X_{n}$ be independent random variables. Write $X=(X_{1},X_{2},\ldots,X_{n})$. Suppose $\psi:F^{n}\to\mathbb{R}$ belongs to $L^{2}$, i.e. $\psi$ is a measurable function such that $E\psi^{2}(X_{1},X_{2},\ldots,X_{n})<\infty$. For each $j\in[n]$, define $E_{j}\psi$ by
@@
In particular, for $\psi\in\mathcal{H}_{t}$, $\psi$ does not depend on $x_{j}$ for $j\notin t$. Then, $(\mathcal{H}_{t})_{t\subset[n]}$ are orthogonal. Denote by $\overline{E}_{t}$ the orthogonal projection from $L^{2}$ onto $\mathcal{H}_{t}$. In fact, $\overline{E}_{t}=\prod_{j\in t}(I-E_{j})\prod_{k\notin t}E_{k}$, where $I$ is the identity map. In particular, $\overline{E}_{\emptyset}$ equals to the usual expectation $E$. Then, we have the orthogonal decomposition
@@
If $\psi$ depends only on $(x_{j})_{j\in s}$ for some $s\subset[n]$ and $t$ is not a subset of $s$, then $\overline{E}_{t}\psi=0$ by definition of $\overline{E}_{t}$ and the fact that $E_{j}\psi=\psi$ for $j\in t\setminus s$. Hence, if $\psi$ depends only on $(x_{j})_{j\in s}$, then we have that
@@
Sometimes, we encounter the random variable $\overline{E}_{t}\psi(X)$. Since $\overline{E}_{t}\psi$ depends only on $(x_{j})_{j\in t}$, we may write $\overline{E}_{t}\psi(X_{t})$ instead of $\overline{E}_{t}\psi(X)$, where $X_{t}$ means the sub-vector $(X_{j})_{j\in t}$.
Then, we are ready for the proof of Theorem 1.6. For a vector $x=(x_{1},x_{2},\ldots,x_{n})$ and a subset $s\subset[n]$, we denote by $x_{s}$ the sub-vector $(x_{j})_{j\in s}$. In this way, we have $X_{S}=(X_{j})_{j\in S}$ and $X_{T}=(X_{j})_{j\in T}$. Since $\partial$ is out of $(F,\mathcal{F})$, we have that
@@
Consider two functions $\varphi$ and $\psi$ such that $E\varphi(S,X_{S})=E\psi(T,X_{T})=0$, $\operatorname{Var}\varphi(S,X_{S})<\infty$ and $\operatorname{Var}\varphi(T,X_{T})<\infty$. For $s\subset[n]$, define a function $\varphi_{s}$ via
@@
Note that $\varphi_{s}$ depends only on $(x_{j})_{j\in s}$. Similarly, we define $\psi_{t}$ for $t\subset[n]$. Then, by independence between $(S,T)$ and $X$, we have that
$\operatorname{Var}\varphi(S,X_{S})$ $=E\varphi^{2}(S,X_{S})$
$=\sum_{s\subset[n]}P(S=s)E\varphi_{s}^{2}(X_{s})$
$=\sum_{u\subset[n]}\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u}),$ (36)
@@
$\varphi_{s,u}=\overline{E}_{u}\varphi_{s}$
and that the last equality of (5.2) comes from the orthogonality in the ANOVA decomposition $\varphi_{s}=\sum_{u:u\subset s}\overline{E}_{u}\varphi_{s}$. Similarly, we have that
$\operatorname{Var}\psi(T,X_{T})=\sum_{u\subset[n]}\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u}),$ (37)
@@
$\leq\sum_{u\subset[n]}\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E\varphi_{s,u}^{2}(X_{u})}\sqrt{E\psi_{t,u}^{2}(X_{u})}.$ (38)
@@
$\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E\varphi_{s,u}^{2}(X_{u})}\sqrt{E\psi_{t,u}^{2}(X_{u})}\\ \leq r_{u}\sqrt{\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})},$ (39)
@@
For $u=\emptyset$, $\varphi_{s,u}=E\varphi_{s}(X_{s})$ and $\psi_{t,u}=E\psi_{t}(X_{t})$. Since $E\varphi(S,X_{S})=E\psi(T,X_{T})=0$, we have that
$\sum_{s\subset[n]}P(S=s)E\varphi_{s}(X_{s})=\sum_{t\subset[n]}P(T=t)E\varphi_{t}(X_{t})=0.$
Define $f(s)=E\varphi_{s}(X_{s})$ and $g(t)=E\varphi_{t}(X_{t})$. Then, we have that $Ef(S)=Eg(T)=0$, $\operatorname{Var}f(S)<\infty$ and $\operatorname{Var}g(T)<\infty$. By definition of maximal correlation coefficients, we have that
$\operatorname{Cov}(f(S),g(T))\leq R(S,T)\sqrt{\operatorname{Var}f(S)}\sqrt{\operatorname{Var}g(T)}.$

### v3

@@ 4 Two-dimensional Lévy processes and bivariate stable distributions
Since $(X_{1},Y_{1})$ is stable, there exists $(c_{n},d_{n})\in\mathbb{R}$ such that $(S_{n}-c_{n},T_{n}-d_{n})$ has the same distribution as $n^{1/\alpha}(X_{1},Y_{1})$. Define
@@
Then $R(X^{(n)},Y^{(n)})=R(S,T)=R(X_{1},Y_{1})$. Moreover, by [Kal21, Theorem 23.14], the process $(X^{(n)}_{t},Y^{(n)}_{t})_{t\geq 0}$ converges towards $(X_{t},Y_{t})_{t\geq 0}$ in the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$. Hence, by Lemma 2.3, we have that
@@
[…]
@@ 4.3 Examples
For $\alpha=2$, $X$ and $Z$ must be Gaussian, and the result goes back to Lancaster [Lan57]. The main contribution is in the case that $0<\alpha<2$, $\lambda\geq 0$, and $X$ are $\alpha$-stable. The result for $\lambda<0$ and symmetric $\alpha$-stable random variables can easily be deduced from (27) by taking Z-Z instead of $Z$. The restriction to positive $\lambda$ in (27) has a reason. In general, the expression for $\lambda<0$ is different. Indeed, we have the following result:
@@
Let $X$ and $Z$ be independent copies of $\alpha$-stable random variables with $0<\alpha<2$. Then the Lévy measure $\nu_{X}(x)\mathrm{d}x$ of $X$ has the following form:
@@
By using characteristic functions and independence between $X$ and $Z$, we find that $(X,Y)=(X,X+\lambda Z)$ is an $\alpha$-stable random vector with Lévy measure $\nu(\mathrm{d}x,\mathrm{d}y)$, where
@@
Recall the notation in the statement of Theorem 1.3. In our case, we have that
@@
$R(X,Y)$ $=\left\|\begin{pmatrix}\sqrt{\frac{c_{+}}{c_{+}+c_{-}|\lambda|^{\alpha}}}&\quad 0\\ 0&\quad\sqrt{\frac{c_{-}}{c_{-}+c_{+}|\lambda|^{\alpha}}}\end{pmatrix}\right\|_{2}$
@@
Secondly, we consider compound Poisson processes.
@@
$\nu(\{(x,y)\in\mathbb{R}^{2}\setminus\{(0,0)\}:x=0\text{ or }y=0\})=0.$
Let $(X_{t},Y_{t})$ be the corresponding Lévy process. Then we have that
@@
The reason is that the number of jumps $N$ of $(X_{t})_{t\geq 0}$ up to time $1$ is equal to that of $(Y_{t})_{t\geq 0}$. Hence, $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq R(N,N)=1$. So, for a generic compound Poisson process, the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ is equal to $1$.
Finally, we show that the maximal correlation coefficient $R(X_{1},Y_{1})$ could be strictly less than $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$.
@@
Let $(B_{t})_{t\geq 0}$ be a (one-dimensional) Brownian motion. Let $(N_{t})_{t\geq 0}$ be a Poisson process of rate $1$ that is independent of $(B_{t})_{t\geq 0}$. Let $X_{t}=B_{t}+N_{t}$. Then $(X_{t},N_{t})_{t\geq 0}$ is a two-dimensional Lévy process. By Theorem 1.2, we have that $R((X_{t})_{t\geq 0},(N_{t})_{t\geq 0})\geq\mathrm{Op}(\nu)=R((N_{t})_{t\geq 0},(N_{t})_{t\geq 0})=1$. Since the maximal correlation coefficient is bounded by $1$ from above, we have that $R((X_{t})_{t\geq 0},(N_{t})_{t\geq 0})=1$. However, $R(X_{1},N_{1})<1$. We prove this by contradiction. Suppose $R(X_{1},N_{1})=1$. Let
$f(x,n)\,\mathrm{d}x=P(X_{1}\in(x,x+\mathrm{d}x),N_{1}=n).$
Then we have that
@@
where $f_{N_{1}|X_{1}}(n|x)=P(N_{1}=n|X_{1}=x)$ and $f_{X_{1}|N_{1}}(x|n)$ is the conditional density of $X_{1}$ given $N_{1}=n$. Here, $C$ is called the mean square contingency of $X_{1}$ and $N_{1}$. We will show that $C<\infty$ as follows: Note that
@@
Then there exists $c>0$ such that for $x\geq e$, we have that
@@
where $\lceil x\rceil$ is the least integer that is greater than or equal to $x$. Then
$f_{X_{1}}(x)=\sum_{n=0}^{\infty}f(x,n)\geq\max_{n\geq 0}f(x,n)=g(x).$
And for $x\geq e$, we have that
@@
Combining (30), (33) and (34), we get (32). Hence, the mean-square contingency $C$ is finite. By [R5́9, Theorem 2], there exist non-degenerate $\varphi(X_{1})$ and $\psi(N_{1})$ such that $\rho(\varphi(X_{1}),\psi(N_{1}))=R(X_{1},N_{1})=1$. Therefore, there exist $c_{1}\neq 0,c_{2}\neq 0$ and $d\in\mathbb{R}$ such that $P(c_{1}\varphi(X_{1})+c_{2}\psi(N_{1})=d)=1$. However, given the value of $N_{1}$, $X_{1}$ could take any value in $\mathbb{R}$. If $P(c_{1}\varphi(X_{1})+c_{2}\psi(N_{1})=d)=1$, then $\varphi$ is constant almost everywhere, which contradicts with the non-degeneracy of $\varphi(X_{1})$. Finally, by contradiction, we prove that $R(X_{1},N_{1})<1$.
For a compound Poisson process, it is still possible that $R(X_{1},Y_{1})<R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$, see the following example:
@@
Let $(M_{t})_{t\geq 0}$ and $(N_{t})_{t\geq 0}$ be two independent Poisson processes with rate $1$. Let $X_{t}=M_{t}-N_{t}$ and $Y_{t}=M_{t}$. Then $(X_{t},Y_{t})_{t\geq 0}$ is a two-dimensional compound Poisson process. We have that $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq\mathrm{Op}(\nu)=1$. Since the maximal correlation coefficient is at most one, we have that $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})=1$. However, we find that $R(X_{1},Y_{1})<1$ by numerical methods, although we do not have a theoretical proof at present. Indeed, let $X^{(n)}_{1}=\min(\max(X_{1},-n),n)$ and $Y^{(n)}_{1}=\min(\max(Y_{1},-n),n)$. Since $X^{(n)}_{1}$ is a measurable function of $X_{1}$ and $Y^{(n)}_{1}$ is a measurable function of $Y_{1}$, we have that $R(X^{(n)}_{1},Y^{(n)}_{1})\leq R(X_{1},Y_{1})$. On the other hand, since $(X^{(n)}_{1},Y^{(n)}_{1})$ converges to $(X_{1},Y_{1})$ in distribution, by Lemma 2.3, we have that $R(X,Y)\leq\liminf_{n\to\infty}R(X^{(n)}_{1},Y^{(n)}_{1})$. Therefore, we must have
@@
Since $(X^{(n)}_{1},Y^{(n)}_{1})$ takes values in a finite set, the maximal correlation $R(X^{(n)}_{1},Y^{(n)}_{1})$ is given by a certain eigenvalue of a finite matrix, which could be found by numerical methods. By numerical calculations, we find that $R(X,Y)=\lim_{n\to\infty}R(X^{(n)}_{1},Y^{(n)}_{1})$ is approximately $0.8321$.
5 Analogs and generalizations of DKS inequality
5.1 Proof of the lower bound in Theorem 1.4
The lower bound
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\geq\sqrt{\max\{P(i\in T):i\in[n]\}}$
is a simple consequence of
$R(X_{i},Y_{i})=\sqrt{P(i\in T)},$
which follows from Lemma 5.1 below.
Let $X$ be a non-degenerate random variable taking values in a general measurable space $(F,\mathcal{F})$. Let $B$ be a Bernoulli random variable independent of $X$. Assume that
P(B=1)=1P(B=0)=p.P(B=1)=1-P(B=0)=p.
Suppose $\partial$ is a special point outside of $F$. Define
@@ 5 Analogs and generalizations of DKS inequality
The maximal correlation coefficient between $X$ and $Y$ is equal to $\sqrt{p}$.
@@
To calculate the maximal correlation coefficient between $X$ and $Y$, we take two functions $f$ and $g$ such that $E[f(X)]=0$, $\operatorname{Var}(f(X))<\infty$, $E[g(Y)]=0$ and $\operatorname{Var}(g(Y))<\infty$. Then we have that
$\operatorname{Var}(g(Y))=E[(g(Y))^{2}]=E[E[(g(Y))^{2}|B]]=pE[(g(X))^{2}]+(1-p)(g(\partial))^{2}.$ (35)
@@
where we use the Cauchy-Schwarz inequality $(E[f(X)g(X)])^{2}\leq E[(f(X))^{2}]E[(g(X))^{2}]$ in the last step. Thus, we have proved that $R(X,Y)\leq\sqrt{p}$.
Finally, there exists $g$ such that
[…]
Take $f=g$. Then
$E[f(X)]=0\text{ and }0<\operatorname{Var}(f(X))=E[(g(X))^{2}]<\infty.$
Moreover, $E[g(Y)]=pE[g(X)]=0$ and $0<\operatorname{Var}(g(Y))<\infty$ by (35). By (5.1), we have that
$R(X,Y)\geq\rho(f(X),g(Y))=\sqrt{p},$
@@
In the above argument, we need the condition that $\partial$ is outside of $F$ to ensure that the distribution of $X$ is non-degenerate conditionally on $X\neq\partial$, which guarantees the existence of $g$ such that $g(\partial)=0$, $E[g(X)|X\neq\partial]=0$ and $0<E[(g(X))^{2}|X\neq\partial]<\infty$. It is possible to allow $P(X=\partial)>0$. However, in such cases, we need to assume that the distribution of $X$ is non-degenerate conditionally on $X\neq\partial$. Otherwise, the formula is no longer correct. For example, we take $B$ to be a Bernoulli random variable with parameter $p=1/2$. Let $X$ be an independent copy of $B$. Take $\partial=0$. Then $Y=BX$. Note that $(X,Y)$ is a bivariate Bernoulli vector. By (12), we find that $R(X,Y)=1/\sqrt{3}<\sqrt{p}=1/\sqrt{2}$.
@@
In this subsection, we prove Theorem 1.5. The key ingredient is the analysis of variance (ANOVA) decomposition developed in [MB07, Appendix I]. For the convenience of the readers, we briefly explain the ANOVA decomposition without proofs. Let $X_{1},X_{2},\ldots,X_{n}$ be independent random variables. Write $X=(X_{1},X_{2},\ldots,X_{n})$. Suppose that $\psi:F^{n}\to\mathbb{R}$ belongs to $L^{2}$, that is, $\psi$ is a measurable function such that $E[\psi^{2}(X_{1},X_{2},\ldots,X_{n})]<\infty$. For each $j\in[n]$, define $E_{j}\psi$ by
@@
In particular, for $\psi\in\mathcal{H}_{t}$, $\psi$ does not depend on $x_{j}$ for $j\notin t$. Then $(\mathcal{H}_{t})_{t\subset[n]}$ are orthogonal. Denote by $\overline{E}_{t}$ the orthogonal projection from $L^{2}$ onto $\mathcal{H}_{t}$. In fact, $\overline{E}_{t}=\prod_{j\in t}(I-E_{j})\prod_{k\notin t}E_{k}$, where $I$ is the identity map. In particular, $\overline{E}_{\emptyset}$ is equal to the usual expectation $E$. Then we have the orthogonal decomposition
@@
If $\psi$ depends solely on $(x_{j})_{j\in s}$ for some $s\subset[n]$ and $t$ is not a subset of $s$, then $\overline{E}_{t}\psi=0$ by the definition of $\overline{E}_{t}$ and the fact that $E_{j}\psi=\psi$ for $j\in t\setminus s$. Hence, if $\psi$ depends solely on $(x_{j})_{j\in s}$, then we have that
@@
Sometimes, we encounter the random variable $\overline{E}_{t}\psi(X)$. Since $\overline{E}_{t}\psi$ depends solely on $(x_{j})_{j\in t}$, we may write $\overline{E}_{t}\psi(X_{t})$ instead of $\overline{E}_{t}\psi(X)$, where $X_{t}$ means the subvector $(X_{j})_{j\in t}$.
We now prove Theorem 1.5. For a vector $x=(x_{1},x_{2},\ldots,x_{n})$ and a subset $s\subset[n]$, we denote by $x_{s}$ the subvector $(x_{j})_{j\in s}$. In this way, we have $X_{S}=(X_{j})_{j\in S}$ and $X_{T}=(X_{j})_{j\in T}$. Since $\partial$ is outside of $(F,\mathcal{F})$, we have the following.
@@
Consider two functions $\varphi$ and $\psi$ such that
$E[\varphi(S,X_{S})]=E[\psi(T,X_{T})]=0,\operatorname{Var}(\varphi(S,X_{S}))<\infty\text{ and }\operatorname{Var}(\varphi(T,X_{T}))<\infty.$
For $s\subset[n]$, define a function $\varphi_{s}$ by
@@
Note that $\varphi_{s}$ depends solely on $(x_{j})_{j\in s}$. Similarly, we define $\psi_{t}$ for $t\subset[n]$. Then by the independence between $(S,T)$ and $X$, we have that
$\operatorname{Var}(\varphi(S,X_{S}))$ $=E[\varphi^{2}(S,X_{S})]$
$=\sum_{s\subset[n]}P(S=s)E[\varphi_{s}^{2}(X_{s})]$
$=\sum_{u\subset[n]}\sum_{s:s\supset u}P(S=s)E[\varphi_{s,u}^{2}(X_{u})],$ (37)
@@
$\varphi_{s,u}:=\overline{E}_{u}\varphi_{s}.$
The last equality of (5.2) comes from the orthogonality in the ANOVA decomposition
$\varphi_{s}=\sum_{u:u\subset s}\overline{E}_{u}\varphi_{s}.$
Similarly, we have that
$\operatorname{Var}(\psi(T,X_{T}))=\sum_{u\subset[n]}\sum_{t:t\supset u}P(T=t)E[\psi_{t,u}^{2}(X_{u})],$ (38)
@@
$\leq$ $\sum_{u\subset[n]}\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E[\varphi_{s,u}^{2}(X_{u})]}\sqrt{E[\psi_{t,u}^{2}(X_{u})]}.$ (39)
@@
$\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E[\varphi_{s,u}^{2}(X_{u})]}\sqrt{E[\psi_{t,u}^{2}(X_{u})]}\\ \leq r_{u}\sqrt{\sum_{s:s\supset u}P(S=s)E[\varphi_{s,u}^{2}(X_{u})]}\sqrt{\sum_{t:t\supset u}P(T=t)E[\psi_{t,u}^{2}(X_{u})]},$ (40)
@@
For $u=\emptyset$, we have that $\varphi_{s,u}=E[\varphi_{s}(X_{s})]$ and $\psi_{t,u}=E[\psi_{t}(X_{t})]$. Since $E[\varphi(S,X_{S})]=0$ and $E[\psi(T,X_{T})]=0$, we have that
$\sum_{s\subset[n]}P(S=s)E[\varphi_{s}(X_{s})]=\sum_{t\subset[n]}P(T=t)E[\varphi_{t}(X_{t})]=0.$
Define $f(s)=E[\varphi_{s}(X_{s})]$ and $g(t)=E[\varphi_{t}(X_{t})]$. Then we have that
$E[f(S)]=E[g(T)]=0,\operatorname{Var}(f(S))<\infty,\operatorname{Var}(g(T))<\infty.$
By the definition of maximal correlation coefficients, we have that
$\operatorname{Cov}(f(S),g(T))\leq R(S,T)\sqrt{\operatorname{Var}(f(S))}\sqrt{\operatorname{Var}(g(T))}.$